THE DOPPLER-FREE TWO-PHOTON ABSORPTION SPECTROSCOPY OF NAPHTHALENE AND THE ZEEMAN EFFECTS

MITSUSHI OKUBO, MASATOSHI MISONO, JINGUO WANG, and \underline{HAJIME KATO}, Molecular Photoscience Research Centre, Kobe University, Nada-ku, Kobe 657-8501, Japan; MASAAKI BABA, Faculty of Integrated Human Studies, Kyoto University, Kyoto 606-8501, Japan.

The Doppler-free spectrum and the Zeeman effect of the A ^{1}B_{1u}(v_{4} = 1 : b_{1u}) ← X ^{1}A_{g}(v = 0) transition of naphthalene in the 33576.4-33578.0 cm^{−1} range has been measured by means of two-photon absorption spectroscopy with counter propagating light beams of identical wavelength within an external cavity. Rotational lines were fully resolved, and 1098 ^{Q(K)}Q(J) lines of J = 0 − 42, K = 0 − 24 were assigned. The molecular constants of the A ^{1}B_{1u}(v_{4} = 1 : b_{1u}) state were determined. Deviations from the line positions calculated using the molecular constants were observed for several lines. Perturbations were not observed for K=0 but were observed to increase with increasing values of K. The perturbations were therefore identified as originating from a parallel Coriolis interaction. The Zeeman splittings for lines of a given J were observed to be maximum at K = 0 and to decrease with increasing K. Analysis of these results indicate that the magnetic moment lies along the c axis (perpendicular to the molecular plane). Any lines broader than our instrumental resolution (5 MHz) were not observed at H = 0 T. The J and K dependence of the Zeeman splittings of the A ^{1}B_{1u}(v_{4} = 1 : b_{1u}) ← X ^{1}A_{g}(v = 0) transitions were observed to be regular. This observation and the small number of perturbed lines, leads to the conclusion that the resonance interaction of the A ^{1}B_{1u} state with the T_{1} ^{3}B_{3u} and T_{2} ^{3}B_{1u} states are small and negligible in the observed region.